A circular playground for a daycare center is to be covered with 9" square tiles. The circle has a diameter of 12 ft. The tiles are packaged in boxes of 25 tiles. How many boxes of tiles are needed?
A. 8 boxes
B. 9 boxes
C. 12 boxes
D. 13 boxes

Answer by stanbon(75887)

A circular playground for a daycare center is to be covered with 9" square tiles. The circle has a diameter of 12 ft. The tiles are packaged in boxes of 25 tiles. How many boxes of tiles are needed?
---------
radius = 6 ft = 72 inches
----------
Area = pi*72^2 = 5184*pi sq in
----------------
Area covered by a box of tiles = 9^2*25 sq in
--------------
# of boxes need = [5184*pi]/[81*25] = 8.0425 boxes
------------------
9 boxes needed
---------------
Cheers,
Stan H.
A. 8 boxes
B. 9 boxes
C. 12 boxes
D. 13 boxes